# Boosting Graph Foundation Model from Structural Perspective

### v1

@@ Abstract
Graph foundation models have recently attracted significant attention due to its strong generalizability. Although existing methods resort to language models to learn unified semantic representations across domains, they disregard the unique structural characteristics of graphs from different domains. To address the problem, in this paper, we boost graph foundation model from structural perspective and propose BooG. The model constructs virtual super nodes to unify structural characteristics of graph data from different domains. Specifically, the super nodes fuse the information of anchor nodes and class labels, where each anchor node captures the information of a node or a graph instance to be classified. Instead of using the raw graph structure, we connect super nodes to all nodes within their neighborhood by virtual edges. This new structure allows for effective information aggregation while unifying cross-domain structural characteristics. Additionally, we propose a novel pre-training objective based on contrastive learning, which learns more expressive representations for graph data and generalizes effectively to different domains and downstream tasks. Experimental results on various datasets and tasks demonstrate the superior performance of BooG. We provide our code and data here: https://anonymous.4open.science/r/BooG-EE42/.
@@ 1 Introduction
Graph learning has achieved outstanding results in a wide range of application fields in recent years, including social networks [22, 13], citation networks [18], molecular graphs [2, 33] and recommendation systems [6]. For graph data containing complex structural information, graph neural networks (GNNs) [25, 11, 29] leverage designed message-passing mechanisms to learn consistent representations for similar instances and apply them to various downstream tasks, such as node classification [11], graph classification [29] and link prediction [31].
Although general GNN models have achieved widespread success, they often face limitations due to training labels and can only perform a single task. In practical applications, accurate ground truth labels often come at a high cost [9]. To address these challenges, graph self-supervised learning (SSL) [26, 30, 9] utilizes the structure or characteristics of the data itself for learning, without the need for explicitly providing labels for the training instances. For example, graph contrastive learning (GCL) [30] supplements supervised signals by comparing the similarity and dissimilarity between different instances in the graph data. Graph SSL methods typically employ a learning pipeline involving “pre-training” and “fine-tuning”. By designing appropriate pre-training tasks, the model is enabled to generalize to downstream tasks after fine-tuning.
@@
To enhance the generalization capability of graph learning for cross-domain graph tasks, graph foundation models based on LLMs aim to learn unified semantic representations on graph data from different domains by text-attribute graphs (TAGs) [14, 28, 7]. However, these approaches are still constrained by the unique structural characteristics of graph data from different domains. For example, citation networks contain citation relationships between academic papers, often exhibiting high levels of clustering, where papers referencing each other may involve similar topics or domains. On the other hand, chemical molecular graphs typically exhibit fixed topological structures and physical-chemical properties. The structural characteristics from different domains make it challenging for models to generalize the learned structural knowledge to other domains. To further enhance the model generalizability, there naturally arises a question: Can we develop a graph foundation model that unifies cross-domain structural characteristics?
To address the aforementioned challenges, in this paper, we propose the Boosting Graph Foundation Model, namely, BooG, which achieves robust performance and generalization capabilities across various graph domains and tasks. First, BooG uses TAGs to map textual descriptions of input data to a unified space by a pre-trained language model (LM). It then introduces virtual super nodes to construct unified structural characteristics for graph data from different domains, addressing the challenge of domain-specific structural characteristics that hinders generalization. Specifically, the virtual super nodes fuse the information of both anchor nodes and class labels, where each anchor node captures the information of a node instance or a graph instance to be classified. Instead of utilizing the raw graph structure, we introduce super nodes and connect them to all the nodes within the neighborhood of an anchor by virtual edges. Details will be illustrated in Figure 3 later. This new linkage allows for effective information aggregation while disregarding the difference in structural connectivity across domains. In this way, cross-domain model generalization can be boosted from structural perspective. Additionally, we propose a novel pre-training objective based on contrastive learning, which learns more expressive representations for graph data and generalizes effectively to different domains and downstream tasks. We conduct extensive experiments on six datasets with diverse domains and tasks. In most cases, our proposed approach leads the state-of-the-art competitors. We also conduct ablation studies to validate the necessity of each component in BooG.
@@ 2 Related Works
Graph Neural Network.
As an effective technique for handling graph-structured data, GNNs [25, 11, 29, 12, 4, 27] can effectively capture complex patterns based on input node features and structural information. For instance, graph attention networks (GAT) [25] utilize attention mechanisms to learn the importance of neighboring nodes and aggregate their information with learned weights. Graph Isomorphism Network (GIN) [29] adopts a unique graph encoding method that emphasizes distinguishing non-isomorphic structures, enhancing the model’s representational capacity. However, the effectiveness of GNNs often relies on a large number of training labels, implying high labeling costs. Additionally, the design of GNNs is often tailored to specific tasks, making it challenging to generalize to other downstream tasks. For instance, GNNs designed for node classification tasks often focus more on local information of the instances, while methods designed for graph classification tasks require more attention to the global information of the instances.
Graph Self-Supervised Learning.
Graph SSL [26, 30, 9] is primarily used to address the problem of lacking labels in graph learning tasks. These methods typically employ learning pipelines that include “pre-training” and “fine-tuning” stages. Through carefully designed contrastive learning or generative pre-training tasks, the model can generalize the learned knowledge to downstream tasks after fine-tuning. For example, GraphCL [30] utilizes the pre-training of graph models by applying a self-supervised contrastive learning task on learned node embeddings. DGI [26] introduces a self-supervised pre-training task aimed at maximizing the mutual information between local node views and global graph views. While graph SSL has been demonstrated to effectively learn instance representations during the pre-training phase, the disparity between pre-training tasks and downstream tasks results in suboptimal performance.
@@
Graph Foundation Model based on LLM.
The advent of foundation models first became a reality in NLP due to the development of LLMs. Foundation models are trained on extensive data and can adapt to a wide range of data and downstream tasks [15]. Existing graph foundation models [14, 7, 28] that integrate LLMs can generally be divided into two categories: (1) Methods with LLMs as the backbone perform various graph tasks by converting graphs into text or tokens for LLMs utilization. (2) Methods with GNN as the backbone typically use pre-trained LLMs to obtain unified representations for instances, enhancing the representation capacity of GNNs and improving generalization. For example, UniGraph [7] proposes a cascade architecture of LM and GNN, along with a self-supervised training objective based on Masked Graph Modeling (MGM), and introduces graph instruction fine-tuning using LLMs to achieve zero-shot prediction capability. OFA [14] utilizes a pre-trained LM to align feature spaces of graphs from different domains, enabling supervised training across all graphs. However, these approaches are still constrained by the structural characteristics of graph data from different domains.
@@ 3.1 Text-Attributed Graphs
3.2 Learning scenarios.
In this paper, we propose a cross-domain and cross-task graph foundation model BooG. This model consists of two parts: a pre-trained encoder and a classifier implemented with an MLP. We first pre-train the encoder and freeze the pre-trained model’s parameters when generalizing to downstream tasks. To train the classifier on downstream tasks, we split the dataset into the train, validation, and test sets, denoted as $\mathcal{D}_{train}$, $\mathcal{D}_{val}$, and $\mathcal{D}_{test}$, respectively. Their label sets are denoted as $\mathcal{Y}_{train}$, $\mathcal{Y}_{val}$, and $\mathcal{Y}_{test}$. To evaluate the generalizability of BooG, we conduct experiments in three different problem settings and adopt three graph-related tasks including node classification, graph classification and link prediction. In supervised learning setting, the classifier will be trained on the training set $\mathcal{D}_{train}$, with the optimal parameters selected based on the validation set $\mathcal{D}_{val}$, and its performance will be finally evaluated on the test set $\mathcal{D}_{test}$. For few-shot learning, $N$-way $K$-shot tasks evaluate the in-context learning ability to apply the pre-learned knowledge to a new domain and task with $N$ classes, each represented by only $K$ labeled examples. For zero-shot learning, $K$ is set to 0, indicating no prior exposure to support instances from the target classes. This setting aims to evaluate a pre-trained model’s ability to generalize and apply its learned knowledge to unseen data categories.
Figure 1: The overall process of BooG. The model’s input includes text attribute graphs and class descriptions. BooG first utilizes a LM to unify different graph data and standardizes the input for node-level and graph-level tasks as sub-graphs. Subsequently, BooG introduces constructed super nodes into sub-graphs and learns to aggregate effective information within the sub-graphs. We freeze the parameters of the pre-trained model and provide the final instance representations through similarity matching. In particular, the similarity matching process can serve as zero-shot learning to predict unseen instances. For supervised learning and few-shot learning scenarios, BooG freezes the parameters of the pre-trained model and generalizes the capabilities of the pre-trained model to multiple downstream tasks by adjusting the parameters of the MLP.
@@ 4 Methodology
We propose a graph foundation model, BooG, which enhances the model generalizability across various graph domains and downstream tasks. In this section, we first introduce how to map the input TAGs from different domains to a unified feature space using a pre-trained language model. Next, we introduce the pre-training algorithm with the goal of learning expressive representations. Finally, we describe how to generalize the pre-trained models to multiple downstream graph tasks in different scenarios. The overall framework of BooG is summarized in Figure 1.
@@
Figure 2: The text format for graph node and class node on Pubmed.
4.1 Unifying Graphs from Diverse Domains
Most GNNs have been designed for specific domains and relied heavily on pre-processed vector features as input, thereby posing challenges in terms of broader applicability and generalization. In this paper, inspired by previous works [14, 32], we introduce TAGs to describe graph nodes using raw texts, as illustrated in Figure 2. Even though the text description comes from different graph domains, we can apply a pre-trained LM to encode these texts into the same semantic space. And in BooG, we adopt Sentence Transformer [21], which has been shown to learn effective sentence embeddings. Let $t_{v}$ denote the text attributes of a graph node $v$, then it can be encoded by the pre-trained LM as follows:
$x(v)=LM(t_{v})\in\mathbb{R}^{d},$ (1)
where $x(v)$ is the output of the LM, and $d$ is the dimension of the output vector. Similarly, let $t_{l}$ be the text attributes of a class node $l$, then it can be encoded by the same pre-trained LM as follows:
$c(l)=LM(t_{l})\in\mathbb{R}^{d}.$ (2)
4.2 Unifying Structure from Diverse Domains
With the advancement of NLP, some works [14, 7] have introduced LMs into graph learning. By mapping graph data from different domains into a unified feature space through TAGs, these efforts aim to enhance the model’s generalization capabilities. However, graph data from different domains often contain unique structural information, posing challenges for existing methods to generalize across different domains. To address the challenges, we propose BooG, which introduces virtual super nodes to construct unified structural space for graph data from different domains.
Sub-graph. Considering that the input graph dataset could consist of a set of nodes for node classification and link prediction tasks, or a set of graphs for graph classification task, we unify the input graph data into sub-graphs composed of anchor nodes and their neighborhoods. Formally, we define a sub-graph $\mathcal{G}_{sub}=\left\{s,\mathcal{N}(s)\right\}$, where $s$ is an anchor node representing a node or a graph instance, and $\mathcal{N}(s)$ denotes its neighborhoods. For node-level tasks, the anchor node $s$ represents an arbitrary node in the input graph, and $\mathcal{N}(s)$ denotes the $k$-hop neighbor nodes of the anchor node $s$. The representation of the anchor node $s$ is the node features encoded by LM, i.e., $h(s)=x(s)$. For graph-level tasks, an anchor node $s$ represents a graph, so the representation of an anchor node $s$ is defined as $h(s)=R_{graph}(\mathcal{G})$, where $R_{graph}$ is a pooling function and the neighborhoods $\mathcal{N}(s)$ include all the nodes in the graph.
Figure 3: The super nodes in sub-graphs.
@@ 4.2 Unifying Structure from Diverse Domains
Super Node. We construct virtual super nodes by fusing each anchor node with all class labels, as shown in Figure 3. For each input node/graph instance, there is only one anchor node $s$ and $C$ super nodes $p$, where $C=\left|\mathcal{L}\right|$ represents the number of classes in a graph dataset. The representation of the $j$-th super node for the $i$-th node/graph can be computed as follows:
$h(p_{ij})=h(s_{i})+\alpha c(l_{j}),j\in\left[1,C\right],$ (3)
where $h(p_{ij})\in\mathbb{R}^{d}$, $h(s_{i})$ represents the representation of an anchor node $s_{i}$, $c(l_{j})$ represents the representation of the $j$-th class label and $\alpha$ is a hyper-parameter. The neighbourhood of a super node $p_{ij}$ is defined as the neighbourhood of its corresponding anchor node $s_{i}$, i.e., $\mathcal{N}(p_{ij})=\mathcal{N}(s_{i})$. We fuse class labels with anchor nodes, aiming to leverage the class labels to assist anchor nodes in identifying and aggregating relevant information within their neighborhoods. We then employ virtual super nodes to construct unified structural characteristics for graph data from different domains.
@@
Encoder. We construct virtual edges $\bar{e}_{pu}$ between each super node $p$ and its neighbor nodes $u\in\mathcal{N}(p)$, and then learn weights for these edges with the attention mechanism. Let the weight $a_{pu}$ denote the importance of the edge $\bar{e}_{pu}$, i.e., the importance of neighbor node $u$ to super node $p$. The weights can be computed as follows:
@@
$a_{pu}=\frac{exp(ReLU(\vec{g}^{T}\left[W_{1}h(p)\parallel W_{2}h(u)\right]))}{{\textstyle\sum_{j\in\mathcal{N}(p)}exp(ReLU(\vec{g}^{T}\left[W_{1}h(p)\parallel W_{2}h(j)\right])}},$ (4)
@@
where $h(u)=x(u)$ represents the node features encoded by the LM, and $W_{1},W_{2}\in\mathbb{R}^{d\times d}$ and $\vec{g}\in\mathbb{R}^{2d}$ are the learnable parameters. Then, we aggregate information from neighborhoods for each super node $p$ to get the corresponding representation. The final representation of super node $p_{ij}$ can be computed as follows:

### v2

# Boosting Cross-Domain and Cross-Task Generalization for Text-Attributed Graphs from Structural Perspective

@@ Abstract
Graph models based on large language models (LLMs) have recently garnered considerable attention due to its significant success. Although existing methods resort to LLMs to learn unified semantic representations across domains, they disregard the unique structural characteristics of graphs from different domains. To address this problem, in this paper, we boost graph models from structural perspective and propose BooG. The model constructs virtual super nodes to unify structural characteristics of graph data from different domains. Specifically, the super nodes fuse the information of anchor nodes and class labels, where each anchor node captures the information of a node or a graph instance to be classified. Instead of using the raw graph structure, the super nodes, along with virtual edges, establish a standardized aggregation mechanism that fuses rich information from neighborhoods and associated class labels, accommodating graph structural characteristics inherent to different domains. Additionally, we propose a novel pre-training objective based on contrastive learning, which learns more expressive representations for graph data and generalizes effectively to different domains and downstream tasks. Experimental results on various datasets and tasks demonstrate the superior performance of BooG. We provide our code and data here: https://github.com/cy623/BooG.
keywords:
Graph learning, Graph foundation model
[…]
@@ 1 Introduction
Graph learning has achieved outstanding results in a wide range of application fields in recent years, including social networks [1, 2], citation networks [3], molecular graphs [4, 5] and recommendation systems [6, 7]. For graph data that encompasses intricate structural information, graph neural networks (GNNs) [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] employ sophisticated message-passing mechanisms to effectively learn consistent representations for instances that share similar characteristics, thereby enabling the application of these representations to a wide range of downstream tasks [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34], such as node classification [9], graph classification [10] and link prediction [35].
Although general GNN models have achieved widespread success, they often face limitations due to training labels and can only perform a single task. In practical applications, accurate ground truth labels often come at a high cost [36]. To address these challenges, graph self-supervised learning (SSL) [37, 38, 36, 39, 40, 41] utilizes the structure or characteristics of the data itself for learning, without the need for explicitly providing labels for the training instances. For example, graph contrastive learning (GCL) [38] supplements supervised signals by comparing the similarity and dissimilarity between different instances in the graph data. Graph SSL methods typically employ a learning pipeline involving “pre-training” and “fine-tuning”. By designing appropriate pre-training tasks, the model is enabled to generalize to downstream tasks after fine-tuning.
@@
To enhance the generalization capability of graph learning for cross-domain graph tasks, graph models based on LLMs aim to learn unified semantic representations on graph data from different domains by text-attribute graphs (TAGs) [46, 47, 48]. However, these approaches are still constrained by the unique structural characteristics of graph data from different domains. For example, citation networks contain citation relationships between academic papers, often exhibiting high levels of clustering, where papers referencing each other may involve similar topics or domains. On the other hand, chemical molecular graphs typically exhibit fixed topological structures and physical-chemical properties. The structural characteristics from different domains make it challenging for models to generalize the learned structural knowledge to other domains.
To address the aforementioned challenges, in this paper, we propose BooG, a pre-trained graph model designed for Boosting cross-domain and cross-task generalization for text-attributed Graphs. BooG unifies the attributes and structures of graph data across different domains, thereby enhancing the capabilities of the pre-trained model to downstream tasks and data from various domains. Specifically, BooG first leverages a pre-trained language model (LM) to encode the attributes of the graph data and class labels into a unified representation space. Next, BooG unifies graph structures by introducing sub-graphs and super nodes. Sub-graphs offer a consistent task template across all levels of tasks, consisting of anchor nodes and their neighbor nodes. The super nodes, along with virtual edges, establish a standardized aggregation mechanism that fuses rich information from neighborhoods and associated class labels, accommodating graph structural characteristics inherent to different domains. Additionally, we propose a novel pre-training objective based on contrastive learning, which enables the model to learn more expressive nodes/graphs representations that generalize effectively to different downstream domains and tasks. Our extensive experiments on seven datasets, covering a range of domains and tasks, show that BooG outperforms state-of-the-art competitors in most cases. We also conduct ablation studies to validate the necessity of each component in BooG.
@@ 2 Related Works
2.1 Graph Neural Network.
As an effective technique for handling graph-structured data, GNNs [8, 9, 10, 49, 50, 51, 52, 53, 54, 55, 56, 26, 57] can effectively capture complex patterns based on input node features and structural information. For instance, graph attention networks (GAT) [8] utilize attention mechanisms to learn the importance of neighboring nodes and aggregate their information with learned weights. Graph Isomorphism Network (GIN) [10] adopts a unique graph encoding method that emphasizes distinguishing non-isomorphic structures, enhancing the model’s representational capacity. However, the effectiveness of GNNs often relies on a large number of training labels, implying high labeling costs. Additionally, the design of GNNs is often tailored to specific tasks, making it challenging to generalize to other downstream tasks. For instance, GNNs designed for node classification tasks often focus more on local information of the instances, while methods designed for graph classification tasks require more attention to the global information of the instances.
2.2 Graph Self-Supervised Learning.
Graph SSL [37, 38, 36, 39, 40, 41] is primarily used to address the problem of lacking labels in graph learning tasks. These methods typically employ learning pipelines that include “pre-training” and “fine-tuning” stages. Through carefully designed contrastive learning or generative pre-training tasks, the model can generalize the learned knowledge to downstream tasks after fine-tuning. For example, GraphCL [38] utilizes the pre-training of graph models by applying a self-supervised contrastive learning task on learned node embeddings. DGI [37] introduces a self-supervised pre-training task aimed at maximizing the mutual information between local node views and global graph views. While graph SSL has been demonstrated to effectively learn instance representations during the pre-training phase, the disparity between pre-training tasks and downstream tasks results in suboptimal performance.
@@
2.3 Graph Foundation Model based on LLM.
The advent of foundation models first became a reality in NLP due to the development of LLMs. Foundation models are trained on extensive data and can adapt to a wide range of data and downstream tasks [58]. Existing graph models [46, 48, 47, 59, 60, 61] that integrate LLMs can generally be divided into two categories: (1) Methods with LLMs as the backbone perform various graph tasks by converting graphs into text or tokens for LLMs utilization. (2) Methods with GNN as the backbone typically use pre-trained LLMs to obtain unified representations for instances, enhancing the representation capacity of GNNs and improving generalization. For example, UniGraph [48] proposes a cascade architecture of LM and GNN, along with a self-supervised training objective based on Masked Graph Modeling (MGM), and introduces graph instruction fine-tuning using LLMs to achieve zero-shot prediction capability. OFA [46] utilizes a pre-trained LM to align feature spaces of graphs from different domains, enabling supervised training across all graphs. However, these approaches are still constrained by structural characteristics of graph from different domains.
@@ 3.1 Text-Attributed Graphs
Figure 1: The overall process of BooG. The model consists of two parts: A pre-trained model based on self-supervised manner and a downstream classifier implemented with a MLP. The model’s input includes text attribute graphs and class descriptions. BooG first utilizes a pre-trained LM to unify different graph data and standardizes the input for node-level and graph-level tasks as sub-graphs. Subsequently, BooG introduces super nodes to establish a standardized aggregation mechanism that fuses rich information from neighborhoods and associated class labels. We freeze the parameters of the pre-trained model and provide the final instance representations through similarity matching. In particular, the similarity matching process can serve as zero-shot learning to predict unseen instances. For supervised learning and few-shot learning scenarios, BooG freezes the parameters of the pre-trained model and generalizes the capabilities of the pre-trained model to multiple downstream tasks by adjusting the parameters of the MLP.
3.2 Message Propagation
Most graph neural networks follow the message-passing mechanism, which consists of two main steps: aggregation and update. We take an arbitrary node $v_{i}$ in the $k$-th layer as an example. The first sub-step is to aggregate information from a node’s neighbors, which is given as:
$\hat{h}_{i}^{(k)}=\texttt{AGGREGATE}(h_{j}^{(k-1)},\forall v_{j}\in\mathcal{N}_{i}).$ (1)
After that, the second sub-step is to update node embeddings:
$h_{i}^{(k)}=\texttt{UPDATE}(h_{i}^{(k-1)},\hat{h}_{i}^{(k)}).$ (2)
In this paper, we implement the encoder by learning information aggregation on the newly constructed structure.
3.3 Learning Scenarios
In this paper, we propose a cross-domain and cross-task graph model BooG. This model consists of two parts: A pre-trained model based on self-supervised manner and a downstream classifier implemented with a Multi-Layer Perceptron (MLP). The pre-trained model, denoted as $f_{\theta}$, operates on a TAG and generates an embedding for each sample (a node or a graph). We freeze the pre-trained model’s parameters when generalizing to downstream tasks. Our model’s generalization ability is evaluated across domains using unseen datasets through three different machine learning problems, and we adopt three graph-related tasks, including node classification, graph classification, and link prediction. We split the dataset into the train, validation, and test sets, denoted as $\mathcal{D}_{train}$, $\mathcal{D}_{val}$, and $\mathcal{D}_{test}$, respectively. Their label sets are denoted as $\mathcal{Y}_{train}$, $\mathcal{Y}_{val}$, and $\mathcal{Y}_{test}$. In supervised learning setting, the classifier will be trained on the training set $\mathcal{D}_{train}$, with the optimal parameters selected based on the validation set $\mathcal{D}_{val}$, and its performance will be finally evaluated on the test set $\mathcal{D}_{test}$. For few-shot learning, $N$-way $K$-shot tasks evaluate the in-context learning ability to apply the pre-learned knowledge to a new domain and task with $N$ classes, each represented by only $K$ labeled examples. For zero-shot learning, $K$ is set to 0, indicating no prior exposure to support instances from the target classes. This setting aims to evaluate a pre-trained model’s ability to generalize and apply its learned knowledge to unseen data categories.
@@ 4 Methodology
In this section, we present the BooG framework, as illustrated in Figure 1.
@@
Figure 2: The text format for graph node and class label on Pubmed.
4.1 Unifying Graph Attributes
Most learned GNNs are tailored to specific domains and rely on pre-processed vector features, which limits their ability to generalize across domains due to the inherently distinct and domain-specific nature of training graph features. Fortunately, since graph attributes can be expressed in natural language [46], in this paper, we unify graph data attributes from different domains by converting input graphs into TAGs and encoding their semantic information using a pre-trained LM. This transformation enables graph nodes to be represented as raw text, as illustrated in Figure 2. Specifically, we employ the Sentence Transformer [62] as our pre-trained LM, known for its capability to generate effective sentence embeddings. Let $t_{v}$ denote the text attributes of a graph node $v$, then it can be encoded by the pre-trained LM as follows:
$x(v)=\text{LM}(t_{v})\in\mathbb{R}^{d},$ (3)
where $x(v)$ is the output of the LM, and $d$ is the dimension of the output vector.
In addition, each class $l\in\mathcal{L}$ of each dataset is associated with a text attribute $t_{l}\in\mathcal{T_{L}}$. The text attributes $t_{l}$ are also encoded by the same pre-trained LM as follows:
$c(l)=\text{LM}(t_{l})\in\mathbb{R}^{d}.$ (4)
It is worth noting that during BooG’s pre-training phase, we rely solely on the textual descriptions of the dataset classes, rather than the class labels of nodes or graphs, thereby establishing a self-supervised learning paradigm, which will be detailed in the following sections.
4.2 Unifying Graph Structures
Graph datasets from different domains exhibit diverse structural characteristics. For example, citation networks often show high clustering coefficients, whereas molecular graphs have fixed, small-world topologies. Moreover, downstream tasks on these datasets can be at different granularities, such as node classification, link prediction, or graph classification. These discrepancies hinder a unified representation learning framework.
Sub-graph. To accommodate varied domains and tasks under a unified framework, we first convert each input graph into a sub-graph centering on an anchor node. Formally, for a given anchor node $s$, we construct a sub-graph $G_{\text{sub}}=\{s,\mathcal{N}(s)\}$, where $\mathcal{N}(s)$ denotes the neighborhood nodes of $s$.
  • For node/edge-level tasks, $s$ is a node from the input graph; $\mathcal{N}(s)$ comprises its $k$-hop neighbors.
  • For graph-level tasks, $s$ represents the whole graph through a pooled embedding $h(s)=R_{\text{graph}}(G)$, and $\mathcal{N}(s)$ contains all nodes in the graph.
This step ensures that all tasks are reduced to a common format centered around an anchor node and its context.
Figure 3: Unifying graph structures by super nodes.
Super Nodes. We propose super nodes to explicitly incorporate class label information into the graph structure and to align structure learning across different domains and tasks.
@@ 4.2 Unifying Graph Structures
For each anchor node $s_{i}$ in a sub-graph, we create $C$ super nodes, where $C=|\mathcal{L}|$ is the number of class labels in the dataset. The $j$-th super node $p_{ij}$ is a virtual node representing the anchor node $s_{i}$ fused with the class label $l_{j}$. The representation is computed as:
$h(p_{ij})=h(s_{i})+\alpha\cdot c(l_{j}),$ (5)
where $h(s_{i})\in\mathbb{R}^{d}$ is the embedding of anchor node $s_{i}$ obtained via the language model, $c(l_{j})\in\mathbb{R}^{d}$ is the embedding of the $j$-th class label, and $\alpha$ is a hyperparameter controlling the weight of class information. The neighborhood of $p_{ij}$ is defined to be the same as that of $s_{i}$, i.e., $\mathcal{N}(p_{ij})=\mathcal{N}(s_{i})$. This design allows the super node to aggregate from the same context as the anchor node, while being explicitly tied to a specific class label.
Anchor Node $s_{i}$ captures instance-specific information. Super Node $p_{ij}$ hypothesizes “what if $s_{i}$ belongs to class $l_{j}$”. Multiple Super Nodes provide multiple hypotheses per anchor node, one for each possible class. These super nodes act as explicit class-guided targets during contrastive learning. Aggregating neighborhood information conditioned on class embeddings facilitates more discriminative and generalizable representations.
Unified Structure with Virtual Edges. The introduction of super nodes enables a standardized, super node-centered structure that unifies the aggregation process across tasks and domains. For each anchor node $s$, we insert all of its corresponding super nodes into its neighborhood, as illustrated in Figure 3. The neighborhood of each anchor node is obtained through sub-graph sampling to preserve its original structural information. Super nodes combine class labels with anchor nodes, leveraging these labels to assist anchor nodes in identifying relevant information within their neighborhoods.
To explicitly construct this unified structure, we introduce virtual edges $\bar{e}_{pu}$ connecting each super node $p$ to its neighborhood nodes $u\in\mathcal{N}(p)$. These virtual edges differ fundamentally from the original edges in the graph $G$ in both purpose and construction. Original edges reflect the inherent data topology (e.g., citations or chemical bonds), whereas virtual edges are artificially constructed to support standardized aggregation conditioned on class labels. Specifically, for each super node $p$, $\mathcal{N}(p)$ inherits from the neighborhood of its anchor node $s$, ensuring super nodes access the same neighborhood information.
Original edges propagate node-level information through message passing based on fixed topology. In contrast, virtual edges allow super nodes to selectively aggregate neighborhood information through learnable attention weights, guiding the prediction towards the correct class label. This design ensures that the final representations of super nodes are both instance-aware and class-informed. The virtual edges play a crucial role in BooG’s unified structure design, enabling consistent and comparable aggregation across different datasets and tasks. While original edges maintain the graph’s inherent topology, virtual edges provide a controlled mechanism for super nodes to gather label-relevant neighborhood information, which is essential for BooG’s generalization capability.
@@
Encoder. We then learn to aggregate information within the newly constructed structure using an attention mechanism. Let $a_{pu}$ denotes the weight of the virtual edge $\bar{e}_{pu}$, indicating the importance of neighbor node $u$ to super node $p$. The weights are calculated as follows:
@@
$a_{pu}=\frac{\exp\left(\text{ReLU}(\vec{g}^{\,T}\left[W_{1}h(p)\parallel W_{2}h(u)\right])\right)}{{\textstyle\sum_{j\in\mathcal{N}(p)}\exp\left(\text{ReLU}(\vec{g}^{\,T}\left[W_{1}h(p)\parallel W_{2}h(j)\right]\right)}},$ (6)
@@
where $h(u)=x(u)$ represents the node features encoded by the LM, $W_{1},W_{2}\in\mathbb{R}^{d\times d}$ and $\vec{g}\in\mathbb{R}^{2d}$ are learnable parameters, and $\parallel$ denotes concatenation. Next, we aggregate information for each super node $p_{ij}$ from its neighborhood to obtain its final representation $\hat{h}(p_{ij})$, computed as: